# Sliding window-aided ordered statistics decoding for short LDPC codes

### v1

@@ Abstract
This paper introduces an innovative approach to the design of efficient decoders that meet the rigorous requirements of modern communication systems, particularly in terms of ultra-reliability and low latency. We enhance an established hybrid decoding framework by proposing an ordered statistical decoding scheme augmented with a sliding window technique. This novel component replaces a key element of the current architecture, significantly reducing average complexity. A critical aspect of our scheme is the integration of a pre-trained neural network model that dynamically determines the progression or halt of the sliding window process. Furthermore, we present a user-defined soft margin mechanism that adeptly balances the trade-off between decoding accuracy and complexity. Empirical results, supported by a thorough complexity analysis, demonstrate that the proposed scheme holds a competitive advantage over existing state-of-the-art decoders, notably in addressing the decoding failures prevalent in neural min-sum decoders. Additionally, our research uncovers that short LDPC codes can deliver performance comparable to that of short classical linear codes within the critical waterfall region of the SNR, highlighting their potential for practical applications.
@@ I Introduction
In the domain of short LDPC codes, addressing decoding failures in BP due to dominant absorbing sets, a combination of recurrent neural network (RNN) models trained for each dominant absorbing set and OSD has claimed near-ML decoding performance [21]. However, the extensive training and complex implementation have limited its adoption. To overcome these challenges, we introduced the NMS-DIA-OSD framework for LDPC codes in prior research [22], which achieves a favorable balance between performance, complexity, and throughput. Specifically, to tackle NMS decoding failures, iterative trajectories are synthesized via a neural model to produce a new reliability metric, decoding information aggregation (DIA). This DIA has been shown to significantly benefit the subsequent adaptive OSD, which incorporates the historical distribution of error patterns into the current decoding task.
Building on the case-by-case strategy of OSD variants, we propose a new sliding window-aided OSD as an extension of our previous work, substituting for the adaptive OSD component in the NMS-DIA-OSD framework. This mechanism leverages a two-layered neural model, with softmax outputs providing soft information for the decision to terminate the current OSD early or continue. Retaining the low latency characteristic of the NMS-DIA-OSD framework, the new framework further reduces computational complexity and demonstrates competitive performance against SOTA decoders for short LDPC codes, offering a better trade-off when handling NMS decoding failures. Additionally, it is observed that short LDPC codes and short eBCH codes of the same rate and length exhibit near-competitive efficacy in decoding performance and computational complexity.
@@ II-B2 DIA and Adaptive OSD
In our prior work [22], we introduced DIA to serve as a substitute for the magnitude of the received sequence in terms of reliability metrics. Specifically, DIA is implemented as a four-layered neural network model that processes the bit trajectory associated with each failed NMS iterative decoding attempt. The output of this model has been statistically validated to enhance reliability measurements, thereby maximizing the effectiveness of OSD of a given order-$p$ for NMS decoding failures in short LDPC codes.
@@
Unlike conventional OSD, the adaptive OSD includes an additional query phase to gather the statistical distribution of actual error patterns associated with existing NMS decoding failures. First, all TEPs with a given order-$p$ are categorized into blocks known as order patterns. Then these order patterns are prioritized based on the ratio of the number of included genuine error patterns to the size of each order pattern. During the OSD decoding phase, the TEPs along the decoding path, which is a sequence of order patterns, are examined when selecting codeword candidates for the optimal estimate of the transmitted codeword. Consequently, steps 4-9 of Algorithm 1 are adapted accordingly. For a more detailed explanation of DIA and adaptive OSD, interested readers are directed to [22].
@@ III-B3 Training Logistics
For each NMS decoding failure, the DIA output generates a distinct minimum distance list for each order pattern within the given decoding path of length $\hat{l}_{th}$. A sequence of $W_{d}$ sliding windows with a unit stride is recorded, forming the training samples. These samples are labeled ’1’ if the global minimum element of the minimum distance list falls within the current window, or ’0’ to signal the continuation of window sliding. After compiling and shuffling a sufficient number of samples, the training dataset is constructed. To counteract class imbalance, we employ APIs in Tensorflow to assign varying weights to individual samples during training.
@@ III-C OSDSW algorithm
4:Initialize the minimum weighted distance $d_{w}$ to $\infty$, and align the initial window with the first $W_{d}$ order patterns.
@@ IV-A Parameter Settings of Codes
For the LDPC CCSDS (128,64) code [10] within the NMS-DIA-OSDSW framework, we set the maximum number of iterations to $T=12$ and the normalization parameter to $\alpha=0.78$, trained at SNR=2.7dB for the NMS component. The DIA model is presumed to be well-trained, and the OSDSW decoding path is predetermined with $Q=6$ segments when partitioning the MRB. The D-O($\lambda_{m},\hat{l}_{th},s_{m}$), which denotes the combination of DIA and OSWSW with $W_{d}=5$, is referenced in the subsequent plots. Here, $\lambda_{m}$ represents the Hamming weight sum threshold for order patterns [22], analogous to the order-$p$ in the original OSD. As benchmarks, PB-OSD [13] and FS-OSD [14], both with order-$p=3$, are SOTA decoders that balance the number of TEPs required against decoding performance for eBCH codes. For PB-OSD, two evaluations are adapted for the two thresholds to assess their effectiveness on LDPC codes, noted as PB0,2 [22]. For FS-OSD, referred to as FS in the legend, parameters are set as $\tau_{E}=6$, $\beta=0.1$, and $\tau_{PSC}=30$ to match the performance of competing decoders.
@@ IV-B Decoding Performance
We begin by assessing the performance of the PB and FS variants for LDPC codes after necessary adaptation. As noted in [23], the NMS scheme is favored for its simplicity, high throughput, and independence from channel noise variance. Thus, we compare the PB and FS variants as post-processors for NMS decoding failures against the proposed OSDSW.
@@
Figures 4 and 5 present the decoding performance in terms of Frame Error Rate (FER) and complexity in terms of $n_{at}$, respectively. PB0, which adheres to the original threshold evaluation from [13], exhibits the poorest FER performance. After re-evaluation, PB2 shows significant improvement with a slight increase in complexity. FS-OSD offers marginally better performance than PB2 but at a considerably higher complexity. In contrast, D-O(3,30,1.0), equivalent to the adaptive OSD with $\lambda_{m}=3$ and $\hat{l}_{th}=30$, is less attractive considering its performance and complexity. However, by extending $\hat{l}_{th}$ to 40 and 50 and reducing $s_{m}$ to 0.9, our sliding window technique effectively manages complexity while outperforming PB2 across most SNR regions of interest.
@@
In Fig. 6, the final FER of our hybrid framework, a product of NMS’ FER and OSDSW’s FER, is compared with existing NBP variants. The NBP-D(10,4,4) [11], which incorporates a learned decimation step, significantly outperforms NMS or BP but falls behind N-D-O(3,60,0.9) by at least 0.5dB at FER=$10^{-3}$. N-D-O(3,60,0.9) is slightly behind the SOTA D10-O-0(25) [21] by about 0.1dB, although its complexity is considerably lower. Both schemes are within 0.3dB of the ML performance, as depicted in Fig. 6.
@@ V Conclusions and Future Research
In this paper, we have introduced a novel sliding window adaptation of the OSD component within the NMS-DIA-OSD framework [22], which maintains the benefits of superior decoding performance, high throughput, and channel knowledge independence. This adaptation is particularly noteworthy for its ability to customize the decoding scheme for each received sequence, significantly reducing complexity. Specifically, the sliding window elements for a given sequence are processed through a lightweight neural model to determine whether to terminate or continue decoding early. The soft margin between the associated probabilities of this binary decision is easily adjustable, allowing for a convenient trade-off between performance and complexity. Our proposed scheme has been shown to outperform SOTA decoders designed for eBCH codes, particularly in the postprocessing of NMS decoding failures for LDPC codes. Furthermore, a comparative analysis between LDPC and eBCH decoders of the same length and rate revealed that LDPC short codes are indeed competitive with eBCH codes, especially in the waterfall SNR region.

### v2

@@ Abstract
This paper introduces an innovative approach to the design of efficient decoders that meet the rigorous requirements of modern communication systems, particularly in terms of ultra-reliability and low-latency. We enhance an established hybrid decoding framework by proposing an ordered statistical decoding scheme augmented with a sliding window technique. This novel component replaces a key element of the current architecture, significantly reducing average complexity. A critical aspect of our scheme is the integration of a pre-trained neural network model that dynamically determines the progression or halt of the sliding window process. Furthermore, we present a user-defined soft margin mechanism that adeptly balances the trade-off between decoding accuracy and complexity. Empirical results, supported by a thorough complexity analysis, demonstrate that the proposed scheme holds a competitive advantage over existing state-of-the-art decoders, notably in addressing the decoding errors prevalent in neural min-sum decoders. Additionally, our research uncovers that short LDPC codes can deliver performance comparable to that of short classical linear codes within the critical waterfall region of the SNR, highlighting their potential for practical applications.
@@ I Introduction
In the domain of short LDPC codes, addressing decoding errors in BP due to dominant absorbing sets, a combination of recurrent neural network (RNN) models trained for each dominant absorbing set and OSD has claimed near-ML decoding performance [21]. However, the extensive training and complex implementation have limited its adoption. To overcome these challenges, we introduced the NMS-DIA-OSD framework for LDPC codes in prior research [22], which achieves a favorable balance between performance, complexity, and throughput. Specifically, to tackle NMS decoding errors, iterative trajectories are synthesized via a neural model to produce a new reliability metric, decoding information aggregation (DIA). This DIA has been shown to significantly benefit the subsequent adaptive OSD, which incorporates the historical distribution of error patterns into the current decoding task.
Building on the case-by-case strategy of OSD variants, we propose a new sliding window-aided OSD as an extension of our previous work, substituting for the adaptive OSD component in the NMS-DIA-OSD framework. This mechanism leverages a two-layered neural model, with softmax outputs providing soft information for the decision to terminate the current OSD early or continue. Retaining the low latency characteristic of the NMS-DIA-OSD framework, the new framework further reduces computational complexity and demonstrates competitive performance against SOTA decoders for short LDPC codes, offering a better trade-off when handling NMS decoding errors. Additionally, it is observed that short LDPC codes and short eBCH codes of the same rate and length exhibit near-competitive efficacy in decoding performance and computational complexity.
@@ II-B2 DIA and Adaptive OSD
In our prior work [22], we introduced DIA to serve as a substitute for the magnitude of the received sequence in terms of reliability metrics. Specifically, DIA is implemented as a four-layered neural network model that processes the bit trajectory associated with each failed NMS iterative decoding attempt. The output of this model has been statistically validated to enhance reliability measurements, thereby maximizing the effectiveness of OSD of a given order-$p$ for NMS decoding errors in short LDPC codes.
@@
Unlike conventional OSD, the adaptive OSD includes an additional query phase to gather the statistical distribution of actual error patterns associated with existing NMS decoding errors. First, all TEPs with a given order-$p$ are categorized into blocks known as order patterns. Then these order patterns are prioritized based on the ratio of the number of included genuine error patterns to the size of each order pattern. During the OSD decoding phase, the TEPs along the decoding path, which is a sequence of order patterns, are examined when selecting codeword candidates for the optimal estimate of the transmitted codeword. Consequently, steps 4-9 of Algorithm 1 are adapted accordingly. For a more detailed explanation of DIA and adaptive OSD, interested readers are directed to [22].
@@ III-B3 Training Logistics
For each NMS decoding error, the DIA output generates a distinct minimum distance list for each order pattern within the given decoding path of length $\hat{l}_{th}$. A sequence of sliding windows of width $W_{d}$ with a unit stride are recorded, forming the training samples. These samples are labeled ’1’ if the global minimum element of the minimum distance list falls within the current window, or ’0’ to signal the continuation of window sliding. After compiling and shuffling a sufficient number of samples, the training dataset is constructed. To counteract class imbalance, we employ APIs in Tensorflow to assign varying weights to individual samples during training.
@@ III-C Proposed OSD algorithm
4:Initialize the current minimum weighted distance $d_{w}$ to $\infty$, and align the initial window with the first $W_{d}$ order patterns.
@@ IV-A Parameter Settings of Codes
For the LDPC CCSDS (128,64) code [10] within the NMS-DIA-OSDSW framework, we set the maximum number of iterations to $T=12$ and the normalization parameter to $\alpha=0.78$, trained at SNR=2.7dB for the NMS component. The DIA model is presumed to be well-trained, and the OSDSW decoding path is predetermined with $Q=6$ segments when partitioning the MRB. The D-O($\lambda_{m},\hat{l}_{th},s_{m}$), which denotes the combination of DIA and OSDSW with $W_{d}=5$, is referenced in the subsequent plots. Here, $\lambda_{m}$ represents the Hamming weight sum threshold for order patterns [22], analogous to the order-$p$ in the original OSD. As benchmarks, PB-OSD [13] and FS-OSD [14], both with order-$p=3$, are SOTA decoders that balance the number of TEPs required against decoding performance for eBCH codes. For PB-OSD, two evaluations are adapted for the two thresholds to assess their effectiveness on LDPC codes, noted as PB0,2 [22]. For FS-OSD, referred to as FS in the legend, parameters are set as $\tau_{E}=6$, $\beta=0.1$, and $\tau_{PSC}=30$ to match the performance of competing decoders.
@@ IV-B Decoding Performance
We begin by assessing the performance of the PB and FS variants for LDPC codes after necessary adaptation. As noted in [23], the NMS scheme is favored for its simplicity, high throughput, and independence from channel noise variance. Thus, we compare the PB and FS variants as post-processors for NMS decoding errors against the proposed OSDSW.
@@
Figures 4 and 5 present the decoding performance in terms of Frame Error Rate (FER) and complexity in terms of $N_{at}$, respectively. PB0, which adheres to the original threshold evaluation from [13], exhibits the poorest FER performance. After re-evaluation, PB2 shows significant improvement with a slight increase in complexity. FS-OSD offers marginally better performance than PB2 but at a considerably higher complexity. In contrast, D-O(3,30,1.0), equivalent to the adaptive OSD with $\lambda_{m}=3$ and $\hat{l}_{th}=30$, is less attractive considering its performance and complexity. However, by extending $\hat{l}_{th}$ to 40 and 50 and reducing $s_{m}$ to 0.9, our sliding window technique effectively manages complexity while outperforming PB2 in most SNR regions of interest.
@@
In Fig. 6, the final FER of our hybrid framework, a product of NMS’ FER and OSDSW’s FER, is compared with the existing NBP variants. The NBP-D(10,4,4) [11], which incorporates a learned decimation step, significantly outperforms NMS or BP but falls behind N-D-O(3,60,0.9) by at least 0.5dB at FER=$10^{-3}$. N-D-O(3,60,0.9) is slightly behind the SOTA D10-O-0(25) [21] by about 0.1dB, although its complexity is considerably lower. Both schemes are within 0.3dB of the ML performance, as depicted in Fig. 6.
@@ V Conclusions and Future Research
In this paper, we have introduced a novel sliding window adaptation of the OSD component within the NMS-DIA-OSD framework [22], which maintains the benefits of superior decoding performance, high throughput, and channel knowledge independence. This adaptation is particularly noteworthy for its ability to customize the decoding scheme for each received sequence, significantly reducing complexity. Specifically, the sliding window elements for a given sequence are processed through a lightweight neural model to determine whether to terminate or continue decoding early. The soft margin between the associated probabilities of this binary decision is easily adjustable, allowing for a convenient trade-off between performance and complexity. Our proposed scheme has been shown to outperform SOTA decoders designed for eBCH codes, particularly in the post-processing of NMS decoding errors for LDPC codes. Furthermore, a comparative analysis between LDPC and eBCH decoders of the same length and rate revealed that LDPC short codes are indeed competitive with eBCH codes, especially in the waterfall SNR region.